# Eaman Jahani

### v1

@@ F.1 Hypotheses and Results
    Successive Prompt Token Sort Ratio: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused increases in the Token Sort Ratio between successive prompts. When we compare the lowest z-score bracket to the highest z-score bracket, we observe a statistically significant difference. However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket. Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on average successive prompt token sort ratio.
@@
    Successive Prompt ’Contains Previous Prompt’ Dummy: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused increases in the likelihood that the subsequent prompt contains the previous prompt. When we compare the lowest z-score bracket to the highest z-score bracket, we observe a statistically significant difference. However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket. Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on the average probability of a prompt containing the preceding prompt.
@@
    No rescaling: Forthcoming
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts. Those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find comparatively identical resultsin terms of statistical significance. Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Logically, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
@@
    No rescaling: Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
@@
    Successive Prompt ’Contains Previous Prompt’ Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences between any of the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on this measure.
@@
Results:
  • CLIP embedding cosine similarity
    No rescaling: Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
  • DreamSim
    No rescaling: Forthcoming
    Z-score rescaling: We observed statistically significant reductions in the variance ratio of prompt-length, participant-level prompt embedding variance, and the percentage of prompts containing the previous prompt when comparing DALL-E 2 to DALL-E 3. We also found statistically significant reductions in the variance ratio of prompt length and participant-level prompt embedding variance when comparing DALL-E 2 and DALL-E 3 (Revised). Our QTE results comparing performance under DALL-E 2 and DALL-E 3 appear in the main text. The results we have seen when comparing DALL-E 3 (Revised) to DALL-E 2 are qualitatively similar, although the QTEs are attenuated due to the prompt revision. Almost all QTEs when comparing DALL-E 3 and DALL-E 3 (Revised) are positive.
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
@@ Results:
In order to test for differences in prompting performance across participant iterations, we pre-registered and conducted stratified two-sample tests with the 3 treatment arms as our strata. We run this test on each participant’s performance improvement, defined as the performance gap between the initial iteration and the best-performing iteration.
@@
    No rescaling: We do not find evidence for statistically significant relationships between DALL-E 2 versus DALL-E 3 (Verbatim), nor between DALL-E 3 (Verbatim) and DALL-E 3 (Revised). However, we do find a statistically significant difference between DALL-E 2 and DALL-E 3 (Revised), whereby DALL-E 2 users showed greater improvement in cosine similarity compared to DALL-E 3 (Revised) users between their first and best attempts. The latter relationship holds even after adjusting for Benjamini-Hochberg multiple testing.
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
@@
    No rescaling: Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
@@ F.2 Deviations From Pre-registration
  • When conducting our analyses to test H4, we observed that the distribution of the z-scores (our independent variable) did not conform to a normal distribution, displaying extremely large (7.1) or small (-21.9) values. To prevent our findings from being disproportionately influenced by these outliers, particularly in linear models, we excluded observations with absolute z-scores greater than 3, which constituted 2.5% of the user-attempt observations. For robustness, we also conducted the stratification analysis including these outliers and found similar results, with the exception that the Average Treatment Effect (ATE) for performance between the top and middle brackets was only marginally significant.

### v2

@@ F.1 Hypotheses and Results
    Prompt Length: e find statistically significant evidence that increases in the Z-score of a realized image causes the subsequent prompt to be longer (either in terms of words or characters). When we test the differences using the trichotomous variable, the difference between the top Z-score bracket and the bottom bracket is statistically significant, with on average longer prompts in the top bracket. However, we do not observe statistically significant differences between either the top or bottom brackets with the middle bracket. When we estimate a linear model at the user-attempt level, we find a similar relationship that is both positive and statistically significant. A unit increase in the Z-score realized quality of an image causes the next prompt to be longer by about 0.3 words on average. Finally, when we estimate the effect at the user-level, we do not find the Z-score of the first prompt to have a statistically significant effect on the length of the subsequent prompts, when using the trichotomous variable or when fitting a linear model.
@@
    Successive Similarity: We find statistically significant evidence that increases in the Z-score of a realized image causes an increase in successive similarity of the prompt. When we test the differences using the trichotomous variable, the difference between the top Z-score bracket (better than expected images) and both the bottom (worse than expected images) and middle (about expected images) brackets is statistically significant, with the top bracket leading to higher similarity with the previous prompt. However, we do not observe a statistically significant difference between the bottom bracket with the middle bracket. When we estimate a linear model, we find a similar relationship that is both positive and statistically significant. An increase in the realized quality (Z-score) of an image causes the next prompt to be more similar to the previous one. When we estimate the effect at the user-level, we do not find statistically significant effects of the first prompt realized quality on average successive similarity, neither when using the trichotomous variable nor when fitting a linear model.
    Successive Prompt Token Sort Ratio: The results are identical to the case of successive cosine similarity described above. We find statistically significant evidence that increases in the Z-score of the image realized in a previous prompt causes an increase in the Token Sort Ratio between successive prompts, using both the trichotomous variable (top vs bottom and middel brackets) and the linear model. At the user-level, we do not find statistically significant effects of the Z-score of the first prompt on average token sort ratio of successive prompts.
    Successive Prompt ‘Contains Previous Prompt’ Dummy: We find statistically significant evidence that increases in the Z-score of a realized image causes an increase in the likelihood that the subsequent prompt contains the previous one. When we test the differences using the trichotomous variable, the difference between the top Z-score bracket and the bottom bracket is statistically significant, with the top bracket more likely to exactly include the previous prompt. However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket. When we estimate a linear model at the user-attempt level, we find a similar relationship that is both positive and statistically significant. An increase in the realized quality of an image causes the next prompt more likely to include the previous one. Finally, when we estimate the effect at the user-level, we do not find the Z-score of the first prompt to have a statistically significant effect on the average probability of a prompt containing the preceding prompt, when using the trichotomous variable or when fitting a linear model.
@@
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores, as opposed to task-iteration Z-scores). When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks). When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
@@
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores). When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find the same pattern of head-to-head statistical significance results. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks). When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
@@
    Successive Prompt ”Contains Previous Prompt” Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences between any of the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on this measure.
@@
F-test Results (with BH Adjusted p-values):
  • DALL-E 3 (revised) vs. DALL-E 2
    Mean prompt length (words): DALL-E 3 Revised has significantly less variance than DALL-E 2 (ratio 0.5217, $p\leq 10^{-4}$).
    Prompt embedding variance: DALL-E 3 Revised also has significantly less variance than DALL-E 2 (ratio 0.7123, $p=2\times 10^{-4}$).
    Cosine similarity with target image, Z-Score: DALL-E 3 Revised has significantly more variance than DALL-E 2 (ratio 1.255, $p=0.0159$).
    Failed to reject null of no differences in variance: Mean raw DreamSim score vs. target image (p = 0.1333), number of topical transitions (by token sort ratio) (p = 0.3294), mean token sort ratio with previous prompt (p = 0.6244), mean percentile rank of DreamSim score (p = 0.9035), mean cosine similarity to previous prompt (p = 0.8748), mean raw CosineSim score vs. target image (p = 0.8301), mean proportion of prompts containing previous prompt (p = 0.5461), Z-score of DreamSim score (p = 0.4016), number of topical transitions (cosine similarity) (p = 0.3505), and mean percentile rank of CosineSim (p = 0.0795).
  • DALL-E 3 (verbatim) vs. DALL-E 3 (revised)
    Mean percentile rank of CosineSim: DALL-E 3 (verbatim) has significantly less variance than DALL-E 3 (revised) (ratio 0.7347, $p=0.0009$).
    Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) also has significantly less variance (ratio 0.7352, $p=0.0009$).
    Mean percentile rank of DreamSim: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7888, $p=0.0141$).
    Cosine similarity with target image, Z-Score: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7961, $p=0.0159$).
    DreamSim with target image, Z-Score: DALL-E 3 (verbatim) has significantly less variance (ratio 0.8195, $p=0.0377$).
    Number of topical transitions (token sort ratio): DALL-E 3 (verbatim) has significantly more variance (ratio 1.2303, $p=0.0301$).
    Failed to reject null of no differences in variance: Mean raw DreamSim score vs. target image (p = 0.1842), mean raw CosineSim score vs. target image (p = 0.5916), prompt embedding variance (p = 0.9771), mean cosine similarity to previous prompt (p = 0.9035), mean prompt length (words) (p = 0.6244), number of topical transitions (cosine similarity) (p = 0.5916), and mean token sort ratio with previous prompt (p = 0.5206).
  • DALL-E 2 vs. DALL-E 3 (verbatim)
    Mean prompt length (words): DALL-E 3 (verbatim) has significantly less variance than DALL-E 2 (ratio 0.553, $p\leq 10^{-4}$).
    Prompt embedding variance: DALL-E 3 (verbatim) also has significantly less variance (ratio 0.7157, $p=2\times 10^{-4}$).
    Mean raw DreamSim score vs. target image: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7501, $p=0.0015$).
    Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7904, $p=0.0141$).
    Mean percentile rank of DreamSim: DALL-E 3 (verbatim) has significantly less variance (ratio 0.8005, $p=0.0159$).
    Failed to reject null of no differences in variance: Mean percentile rank of CosineSim (p = 0.1842), Z-score of DreamSim score (p = 0.3294), Mean Raw CosineSim (p = 0.8151), Z-score of CosineSim score with target image (p = 0.9906), mean token sort ratio with previous prompt (p = 0.8748), mean cosine similarity to previous prompt (p = 0.807), number of topical transitions (token sort ratio) (p = 0.3505), and number of topical transitions (cosine similarity) (p = 0.085).
QTE Highlighted Results: Through our QTE plots, we see clearest evidence of dispersion being reduced for:
  • Z-score of cosine similarity with respect to the target image (calculated within task-iteration) for the DALLE-3 Revised vs. DALLE-2 comparison.
And we see the clearest evidence of dispersion increasing for:
  • Mean prompt length for all three pairwise model comparisons (DALLE-3 Verbatim vs. DALLE-2, DALLE-3 Revised vs. DALLE-2, and DALLE-3 Verbatim vs. DALLE-3 Revised)
  • Raw Dreamsim performance for DALLE-3 Verbatim vs. DALLE-2 and DALLE-3 Revised vs. DALLE-2 comparisons.
@@
In order to test for differences in prompting performance improvement across participant iterations, we pre-registered and conducted stratified two-sample tests with the 3 treatment arms as our strata. The two samples represent the best and initial scores. We make these comparisons overall and within strata.
@@
    No rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Z-score rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Percentile rank rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
@@
    No rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Z-score rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Percentile rank rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
@@ F.2 Deviations From Pre-registration
  • In our pre-registration, we were not explicit enough on how to compute Z-score measures of performance. As described in section C.2, Z-scores are almost always computed within image-attempt to adjust for variations across target images or attempts. The only exception is with any analysis that compares performance across attempts where Z-scores need to computed within target images. The pre-registration anticipated that Z-scores for any analysis would be computed only within target images.
@@
  • When conducting our analyses to test H4, we observed that the distribution of the Z-scores (our independent variable) did not conform to a normal distribution, displaying extremely large (7.1) or small (-21.9) values. To prevent our findings from being disproportionately influenced by these outliers, particularly in linear models, we excluded observations with absolute Z-scores greater than 3, which constituted 2.65% of the user-attempt observations. For robustness, we also conducted the stratification analysis including these outliers and found almost identical results, with the exception that the difference in performance between the top bracket and middle or bottom brackets was in some cases only marginally significant.
  • In our pre-registration, we had included a rescaled version of our main performance metrics described in B.5.1 based on GPT-4V for robustness check. This outcome variable would have been rescaled by measuring its distance to the outcome variable obtained using a prompt generated by GPT-4V in response to the target image. Given that the quality of the prompt from GPT-4V model is a constant for each target image, this would have amounted to subtracting a constant from the unscaled CLIP cosine similarity or dream sim scores. Furthermore, since all our analysis involving performance as the dependent variable adjust for the target image, either through post-stratification or as a covariate in a linear model, the results from using GPT-4V scaling would have been identical to the the unscaled measures of performance. For this reason, we have not included the robustness checks involving GPT-4V for our pre-registered hypothesis.